OpenStudy (anonymous):

solve y-y1=m(x-x1) for x1.

OpenStudy (dumbcow):

get x1 by itself divide by m on both sides (y-y1)/m = x-x1 subtract x on both sides (y-y1)/m - x = -x1 flip the signs (divide by -1) x - (y-y1)/m = x1